OpenStudy (anonymous):

Please help. Just need answer. Solve for x. Round to nearest hundredth 5=ln(x+1)

OpenStudy (freckles):

write in the equivalent exponent form

OpenStudy (freckles):

\[b=\ln(a) \implies e^{b}=a\]

OpenStudy (anonymous):

Would it be 147.41 ??

OpenStudy (freckles):

so I guess you wrote in the exponential form... \[e^5=x+1 \\ \text{ last step is \to subtract 1 on both sides } \\ e^5-1=x\] well and then you can approximate the number by rounding to nearest hundredth seems good

OpenStudy (anonymous):

So 147.41?

OpenStudy (freckles):

yes

OpenStudy (anonymous):

Yay thank you (: